The freezing point depression of a solution made b
Answer #1

i) depression in freezing point ΔTf = 7.19 oC - (-0.98 oC) = 8.17 oC

ii)      mass of solute = volume x density = 5 ml x 6.18 g/ml = 30.9 g

   molar mass of unknown, M = ?

   mass of solvent in Kg , W = 0.5 kg

molality = (w/M) x (1/W in kg)

                    =  (30.9 g/M) x (1/0.5 kg)        

   = (30.9 /M) x (1/0.5) m

iii)       ΔTf = kf x molality

8.17 oC = 9.8 oC m-1 x (30.9 /M) x (1/0.5) m

M = 74.1 g/mol

Hence,

Molar mass of unknown = 74.1 g/mol